OpenStudy (anonymous):

A survey found that 80% of college students believe their online education courses are as good or superior to courses that use traditional face to face instruction. a. Give the null hypothesis for testing the claim made by the survey. (I know this answer is .80) b. Give the rejection region for a two tailed test conducted at a= .10

OpenStudy (anonymous):

I just need help with b.

OpenStudy (anonymous):

\[\frac{ \alpha }{ 2 }=0.5-\frac{ .01 }{ 2 }=0.45\]

OpenStudy (anonymous):

then to the table u gonna get 1.645

OpenStudy (anonymous):

reject if

OpenStudy (anonymous):

Z<-1.645 or Z>1.645